OpenStudy (anonymous):

A ball moves in a straight line and has an acceleration of a(t) = 2t + 5. Find the position function of the ball if its initial velocity is -3 cm per second and its initial position is 12 cm.

OpenStudy (amistre64):

ok; tell me what you get when you integrate a(t) please

OpenStudy (anonymous):

just t ?

OpenStudy (amistre64):

i think you confusing the issue with derivatives... try to think about what the function would look like that derives down to a(t). in other words: f(x) = ??? such that; f'(x) = 2t +5

OpenStudy (anonymous):

F(t)=2t+5

OpenStudy (amistre64):

lets test that: F(t) = 2t+5 ; when we derive it we get... F'(t) = 2 then right? so that wont work

OpenStudy (amistre64):

think of integrateing as "undoing" a derivative

OpenStudy (amistre64):

the power rule for derivatives is: X^n --> nX^(n-1); in this case n=2 right?

OpenStudy (amistre64):

so it comes from something that resembles t^2 at least. so lets test this option and see where it gets us. t^2 derives to .... 2t. its a perfect match :)

OpenStudy (amistre64):

now we focus on the "5". How do we get a derivative that is 5? its just a constant sitting there without a variable ...

OpenStudy (anonymous):

isn't the variable 12?

OpenStudy (amistre64):

i dont think that 12 can be a variable; many people would confuse it to be the number 12 so that wouldnt be a good option in my book :)

OpenStudy (amistre64):

the constant rule for derivatives says that if K is a constant then: Kx derives to: K; in this case K=5 so lets try 5t as its antiderivative 5t derives to 5, it fits :) this brings us to v(t) = t^2 + 5t; plus some constant that we need to solve for to anchor this floater to the y axis v(t) = t^2 +5t + C

OpenStudy (amistre64):

the term "initial" is code for : when t=0, so we can solve this v(t) specifically by making t=0 and the value of v(0) = -3 as stated in the problem

OpenStudy (amistre64):

-3 = (0)^2 + 5(0) + C -3 = C v(t) = t^2 +5t -3; and in order to get to the position function we antiderive this function

OpenStudy (anonymous):

I hate! and absolutely hate anti derivatives!

OpenStudy (amistre64):

t^2 comes from some for of t^3...but when we derive t^3 we get: [t^3]' = 3t^2 .... we get a pesky 3 in the way, so we need to get rid of it what number can we divide by to get rid of the 3? im gonna say "3" lol (t^3)/3 might work for us, lets test it [(t^3)/3]' = (3/3)t^2 = t^2 .... it works!! :)

OpenStudy (amistre64):

the power rule for antiderivatives is simply: X^n comes from [X^(n+1)]/(n+1) notice the t^2 antiderives to: (t^3)/3 as a result

OpenStudy (anonymous):

I thought we had to find s(t) ?

OpenStudy (amistre64):

5t follows the same rule; the exponent being a "1" \[\frac{5t^{1+1}}{(1+1)}=\frac{5}{2}t^2\]

OpenStudy (amistre64):

we are..we are; but it doesnt just magically appear :)

OpenStudy (anonymous):

So would it be s(t)= 3t^3+5t^2-3t+12 ???

OpenStudy (amistre64):

now for that " -3"; the rule is simply, attach the variable to it. -3 antiderives to -3t ... plain and simple when we put this all together we get v(t) = t^2 +5t -3 antiderives to: s(t) = (t^3)/3 +(5/2)t^2 -3t +C

OpenStudy (anonymous):

I think I narrowed it down to that. Idk if it's right

OpenStudy (anonymous):

amistre help me ; ) plzzz

OpenStudy (anonymous):

Amistre is my baby! leave us alone

OpenStudy (amistre64):

and you got ahead of me but yes :) at the initial condition, code for t=0, we need s(0) = 12 12 = 0+0+0+C 12 = C s(t) = (t^3)/3 +(5/2)t^2 -3t + 12 :)

OpenStudy (amistre64):

youre structure is off a bit i think

OpenStudy (amistre64):

s(t)= 3t^3+5t^2-3t+12 is not s(t) = (t^3)/3 +(5/2)t^2 -3t + 12

OpenStudy (anonymous):

Yeah I found my mistake (:

OpenStudy (anonymous):

omg I get it! kinda ! but I understood it :D

OpenStudy (amistre64):

lol good :)

OpenStudy (anonymous):

babe! now help me pllz!!!